Circular No. 8698

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2362 CYGNI = NOVA CYGNI 2006 H. Yamaoka, Kyushu University, writes that several Japanese observers

took low-resolution spectra around Apr. 5.8 UT of the possible nova reported on IAUC 8697, including K. Kinugasa (Gunma Astronomical Observatory, 1.5-m telescope; range 400–800 nm, resolution ∼ 500); H. Naito and S. Ozaki (Nishi-Harima Astronomical Observatory, 2.0-m NAYUTA telescope; range 420–790 nm, resolution ∼ 1200); and M. Fujii (Ibara, Okayama, Japan; range 390–860 nm). The spectrum shows a rather flat continuum, on which an Hα line shows a clear P-Cyg profile (FWHM of the emission is 500 km/s; the absorption minimum is 700 km/s bluer than the emission peak). Other lines, including Hβ and Hγ, are seen in absorption. This suggests that the object is a classical nova caught at or slightly before its maximum light. Na I D absorption line (EW = 0.3 nm) indicates that some interstellar reddening exists.

N. N. Samus, Russian Academy of Sciences, informs us that this nova has been assigned the designation V2362 Cyg.

Further to IAUC 8697, H. Nishimura provides an additional photographic magnitude of V2362 Cyg: Apr. 5.79 UT, 7.9. Visual magnitude estimates: Apr. 5.427, 8.5: (J. D. West, Mulvane, KS); 5.97, 8.5 (M. Reszelski, Szamotuly-Galowo, Poland).

SUPERNOVAE 2006bk AND 2006bl Two additional apparent supernovae have been reported from unfil-

tered CCD images by T. Boles (cf. IAUC 8684):

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006bk Apr. 3.985 15^{h}04^{m}33.^{s}61 +35^{o}57′51′′.1 16.9 10′′.6 E, 3′′.3 S
2006bl Apr. 4.063 15 39 50.62 +14 11 21.3 17.3 1′′.5 W, 6′′.2 N

Additional magnitudes for 2006bk in MCG +06-33-20: 1994 May 3 UT, [20.5 (Digitized Sky Survey, red); 1998 Mar. 17, [21.0 (DSS, blue); 2006 Jan. 28, [19.5 (Boles); Feb. 7, [19.4 (T. Puckett, Ellijay, GA); Apr. 4.345, 17.4 (Puckett and L. Cox; independent discovery); 4.933, 16.9 (Boles); 5.232, 17.1 (Puckett). Puckett provides the following position end figures for 2006bk: 33^{s}.68, 50′′.6 (offset 11′′.8 east, 3′′.6 south). Additional magnitudes for 2006bl in MCG +02-40-9: 1990 June 14, [21.0 (DSS, blue); 1992 June 20, [20.5 (DSS, red); 2005 June 10, [19.5 (Boles); 2006 Jan. 28, [19.5 (Boles); Apr. 4.943, 17.3 (Boles).

2006 April 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green